Find the general solutions of the following system

  1. dx/dt = -4x+2y , dy/dt = 5x/2 + 2y

2. X' = ( -1 3

- 3 5) X

3. X' = ( 4 -5

5 -4)X



1
Expert's answer
2021-04-29T16:43:18-0400

Consider a general system of linear ODE with two variables.

"\dot{X}(t)=AX(t)", "X\in R^2", "A\in M(2,R)"

The solution can be expressed via exponent of the matrix "(At)" :

"X(t)=e^{At}X(0)", where "e^{At}=\sum\limits_{n=0}^{+\infty}\frac{t^n}{n!}A^n"


2. Let "A=\begin{pmatrix}-1 & 3 \\ -3 & 5\end{pmatrix}" . The characteristic polinomial of A is "\lambda^2-4\lambda+4=(\lambda-2)^2" .

By the Cayley–Hamilton theorem we have "(A-2I)^2=0" . Then

"e^{At}=e^{2tI}e^{(A-2I)t}=e^{2t}\sum\limits_{n=0}^{+\infty}\frac{t^n}{n!}(A-2I)^n=e^{2t}(I+t(A-2I))"

"e^{At}=e^{2t}\begin{pmatrix}1-3t & 3t \\ -3t & 1+3t\end{pmatrix}"

Answer. "X(t)=e^{At}X(0)=e^{2t}\begin{pmatrix}1-3t & 3t \\ -3t & 1+3t\end{pmatrix}X(0)" .


3. Let "A=\begin{pmatrix}4 & -5 \\ 5 & -4\end{pmatrix}". The characteristic polinomial of A is "\lambda^2+9=0."

By the Cayley–Hamilton theorem we have "A^2+9I=0", i.e. "A^2=-9I" . Then

"e^{At}=\sum\limits_{n=0}^{+\infty}\frac{t^{2n}}{(2n)!}A^{2n}+\sum\limits_{n=0}^{+\infty}\frac{t^{2n+1}}{(2n+1)!}A^{2n+1}=\sum\limits_{n=0}^{+\infty}\frac{(-9)^n t^{2n}}{(2n)!}I+\sum\limits_{n=0}^{+\infty}\frac{(-9)^nt^{2n+1}}{(2n+1)!}A=\sum\limits_{n=0}^{+\infty}(-1)^n \frac{(3t)^{2n}}{(2n)!}I+\frac{1}{3}\sum\limits_{n=0}^{+\infty}(-1)^n\frac{(3t)^{2n+1}}{(2n+1)!}A=\cos(3t)I+\frac{1}{3}\sin(3t)A"

"e^{At}=\frac{1}{3}\begin{pmatrix}3\cos 3t +4\sin 3t & -5\sin 3t \\ 5\sin 3t & 3\cos 3t -4\sin 3t\end{pmatrix}"

Answer. "X(t)=e^{At}X(0)=\frac{1}{3}\begin{pmatrix}3\cos 3t +4\sin 3t & -5\sin 3t \\ 5\sin 3t & 3\cos 3t -4\sin 3t\end{pmatrix}X(0)"


1. Assume that the matrix A has two different real roots "\lambda_1" and "\lambda_2" . The characteristic polinomial of A is "\lambda^2-(\lambda_1+\lambda_2)\lambda+\lambda_1\lambda_2=0". By the Cayley–Hamilton theorem we have "A^2-(\lambda_1+\lambda_2)A+\lambda_1\lambda_2I=0". Let "X_1" and "X_2" be the eigenvectors of the matrix A, corresponding to the eigenvalues "\lambda_1" and "\lambda_2" respectively. They form a basis of the vector space "R^2", hence "X(0)=c_1X_1+c_2X_2" with some unknown coefficients "c_1" and "c_2". Then

"e^{At}X_i=\sum\limits_{n=0}^{+\infty}\frac{t^n}{n!}A^nX_i=\sum\limits_{n=0}^{+\infty}\frac{t^n}{n!}\lambda_i^nX_i=e^{\lambda_it}X_i". So

"X(t)=e^{At}X(0)=e^{At}(c_1X_1+c_2X_2)=c_1e^{\lambda_1t}X_1+c_2e^{\lambda_2t}X_2"

Notice that

"(\lambda_1 I-A)X(0)=(\lambda_1 I-A)(c_1 X_1+c_2 X_2)=c_2(\lambda_1-\lambda_2) X_2"

"(\lambda_2 I-A)X(0)=(\lambda_2 I-A)(c_1 X_1+c_2 X_2)=c_1(\lambda_2-\lambda_1) X_1"

therefore,

"c_2X_2=\frac{1}{\lambda_1-\lambda_2}(\lambda_1 I-A)X(0)",

"c_1X_1=\frac{1}{\lambda_2-\lambda_1}(\lambda_2 I-A)X(0)" and

"X(t)=c_1e^{\lambda_1t}X_1+c_2e^{\lambda_2t}X_2=e^{\lambda_1t}\frac{1}{\lambda_1-\lambda_2}(\lambda_1 I-A)X(0)+e^{\lambda_2t}\frac{1}{\lambda_2-\lambda_1}(\lambda_2 I-A)X(0)"

If we take "A=\begin{pmatrix}-4 & 2 \\ 5/2 & 2\end{pmatrix}", then the characteristic polynomial equals to "\lambda^2+2\lambda-13", its roots are "\lambda_1=-1+\sqrt{14}" and "\lambda_2=-1-\sqrt{14}" .

"X(t)=e^{\lambda_1t}\frac{1}{\lambda_1-\lambda_2}(\lambda_1 I-A)X(0)+e^{\lambda_2t}\frac{1}{\lambda_2-\lambda_1}(\lambda_2 I-A)X(0)" implies

"X(t)=\frac{e^{(-1+\sqrt{14})t}}{2\sqrt{14}}((-1+\sqrt{14}) I-A)X(0)-\frac{e^{(-1-\sqrt{14})t}}{2\sqrt{14}}((-1-\sqrt{14}) I-A)X(0)"

"X(t)=\frac{e^{-t}}{\sqrt{14}}((\sqrt{14}\cosh\sqrt{14}t) I-(I+A)\sinh \sqrt{14}t)X(0)"

"X(t)=e^{-t}((\cosh\sqrt{14}t) I-\frac{\sinh \sqrt{14}t}{\sqrt{14}}(I+A))X(0)"

Answer. "X(t)=e^{-t}((\cosh\sqrt{14}t) I-\frac{\sinh \sqrt{14}t}{\sqrt{14}}(I+A))X(0)"